# Inference-Time Alignment for Diffusion Models via Doob’s Matching

### v1

@@ 3.2  Controllable diffusion models with Doob’s $h$-transform
In this subsection, we achieve inference-time alignment by incorporating guidance into the base diffusion models via Doob’s $h$-transform (Rogers and Williams, 2000; Särkkä and Solin, 2019; Heng et al., 2024; Chewi, 2025). We begin by constructing a target path measure $\mathbb{Q}$} on the filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}})$. We assume that $\mathbb{Q}$} is absolutely continuous with respect to the base path measure $\mathbb{P}$} defined in Section (2.2). By the Radon-Nikodym theorem, this relationship is characterized by the existence of an $\mathbb{F}$}-adapted process $(L_{t})_{0\leq t\leq T}$ with $L_{t}\geq 0$, such that for every $t\in[0,T]$:
@@
The process $L_{t}$ is known as the Radon-Nikodym derivative process, representing the likelihood ratio between the target and base measures conditioned on the filtration ${\mathcal{F}}_{t}$.
@@
The rest of our derivation proceeds in two steps. First, we characterize the stochastic dynamics of $\mathbf{X}_{t}^{\leftarrow}$ under the target measure $\mathbb{Q}$}. Second, by invoking the weak uniqueness of solutions to stochastic differential equations (Øksendal, 2003, Lemma 5.3.1), we construct a controllable diffusion process under the base path measure $\mathbb{P}$} whose terminal distribution coincides with the target tilted distribution $q_{0}$.
@@
The proof of Proposition 3.1 is provided in Appendix B. With the dynamics of $\log L_{t}$ established, we derive the stochastic dynamics of $\mathbf{X}_{t}^{\leftarrow}$ under the target measure $\mathbb{Q}$} via Girsanov’s theorem (Øksendal, 2003, Theorem 8.6.6).
@@
Let the base process $\mathbf{X}_{t}^{\leftarrow}$ satisfy the SDE (2.4) under the path measure $\mathbb{P}$}, and let $\mathbb{Q}$} be the target path measure defined by (3.9). Assume that the Novikov condition holds:
@@
where $\mathbf{B}_{t}$ is a Brownian motion under $\mathbb{P}$}. Then $\widetilde{\mathbf{B}}_{t}$ is a standard Brownian motion under $\mathbb{Q}$}. Further, under the path measure $\mathbb{Q}$}, the base process $\mathbf{X}_{t}^{\leftarrow}$ in (2.4) evolves according to:
@@
The proof is deferred to Appendix B. By the construction in (3.9), the law of $\mathbf{X}_{T}^{\leftarrow}$ under $\mathbb{Q}$} coincides with the target tilted distribution $q_{0}$.
@@
Controllable diffusion process under the base measure
@@
To address this, we construct a surrogate process $\mathbf{Z}_{t}^{\leftarrow}$ on the base probability space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$ that adopts the drift derived for $\mathbb{Q}$}:
@@
3.3  Doob’s matching with gradient penalty
Up to now, we have found a controllable diffusion process (3.14) that can generate samples from the target tilted distribution $q_{0}$. However, the Doob’s $h$-function defined as (3.10) is intractable. This subsection aims to estimate the Doob’s guidance $\nabla\log h_{t}^{*}$ in (3.14).
For every $t\in(0,T)$, the Doob’s $h$-function $h_{t}^{*}\coloneq h^{*}(t,\cdot)$ is the unique solution to the following optimization problem:
(3.15) $\min_{h_{t}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}}{\mathbb{E}}^{{\mathbb{P}}}\big[\|h_{t}(\mathbf{X}_{t}^{\leftarrow})-h_{t}^{*}(\mathbf{X}_{t}^{\leftarrow})\|_{2}^{2}\big]=\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}.$
However, since this objective functional includes unknown function $h_{t}^{*}$, we cannot use this explicit Doob’s matching objective functional. Based on the property of the conditional expectation (3.10), we now propose an alternative objective functional, named implicit Doob’s matching:
@@ 3.3  Doob’s matching with gradient penalty
$={\mathbb{E}}_{\mathbf{X}_{0}\sim p_{0}}{\mathbb{E}}_{{\bm{\varepsilon}}\sim{\mathcal{N}}({\mathbf{0}},\mathbf{I}_{d})}\big[\|h_{t}(\mu_{T-t}\mathbf{X}_{0}+\sigma_{T-t}{\bm{\varepsilon}})-w(\mathbf{X}_{0})\|_{2}^{2}\big],$
where $w:{\mathbb{R}}^{d}\to{\mathbb{R}}_{\geq 0}$ is the known weight function defined as (3.1).
The following proposition bridges the objective functionals of the explicit Doob’s matching (3.15) and the implicit Doob’s matching (3.16).
@@
The proof of Proposition 3.4 can be found in Appendix B. Since the expectation in the population risk (3.16) is computationally intractable, we approximate it by empirical risk using independent and identically distributed samples:
(3.17) $\widehat{{\mathcal{J}}}_{t}(h_{t})\coloneq\frac{1}{n}\sum_{i=1}^{n}\|h_{t}(\mu_{T-t}\mathbf{X}_{0}^{i}+\sigma_{T-t}{\bm{\varepsilon}}^{i})-w(\mathbf{X}_{0}^{i})\|_{2}^{2},$
where $\mathbf{X}_{0}^{1},\ldots,\mathbf{X}_{0}^{n}$ are independent and identically distributed random variables drawn from the base distribution $p_{0}$, and ${\bm{\varepsilon}}_{1},\ldots,{\bm{\varepsilon}}_{n}$ are independent standard Gaussian random variables. A naive approach for estimating the Doob’s $h$-function $h_{t}^{*}$ (3.10) involves vanilla empirical risk minimization:
(3.18) $\widehat{h}_{t}\in\operatorname*{arg\,min}_{h_{t}\in{\mathscr{H}}_{t}}\widehat{{\mathcal{J}}}_{t}(h_{t}),$
Limitations of vanilla regression
@@
It is crucial that estimating the Doob’s guidance $\nabla\log h_{t}^{*}$ in (3.14) necessitates not only the estimation of the Doob’s $h$-function $h_{t}^{*}$ but also its gradient $\nabla h_{t}^{*}$, as expressed by
$\nabla\log h_{t}^{*}(\mathbf{z})=\frac{\nabla h_{t}^{*}(\mathbf{z})}{h_{t}^{*}(\mathbf{z})},\quad\mathbf{z}\in{\mathbb{R}}^{d}.$
Previous research, including classifier guidance (Dhariwal and Nichol, 2021), Monte Carlo regression Uehara et al. (2025b, Section 2.2) and Ouyang et al. (2024), approaches this challenge by approximating the Doob’s $h$-function $h_{t}^{*}$ with the vanilla empirical risk minimizer $\widehat{h}_{t}$ (3.18), and subsequently approximating the gradient of the Doob’s $h$-function $\nabla h_{t}^{*}$ via the plug-in gradient estimator $\nabla\widehat{h}_{t}$.
Although rigorous theoretical studies (Bauer and Kohler, 2019; Nakada and Imaizumi, 2020; Schmidt-Hieber, 2020; Kohler and Langer, 2021; Farrell et al., 2021; Kohler et al., 2022; Jiao et al., 2023) demonstrate that the vanilla empirical risk minimizer $\widehat{h}_{t}$ converges to the Doob’s $h$-function $h_{t}^{*}$ in the $L^{2}$-norm, the convergence of the corresponding plug-in gradient estimator $\nabla\widehat{h}_{t}$ cannot be guaranteed, which has been mentioned in Mou (2025, Section 3.2.2). To illustrate this fundamental limitation, consider the sequence of functions $f_{n}:{\mathcal{I}}\rightarrow{\mathbb{R}}$, defined as $x\mapsto n^{-1}\sin(nx)$, and the zero function $f_{0}:{\mathcal{I}}\rightarrow{\mathbb{R}}$, where ${\mathcal{I}}:=[0,2\pi]$. While $\|f_{n}-f_{0}\|_{L^{p}({\mathcal{I}})}\rightarrow 0$ as $n\rightarrow\infty$, we observe that $\lim_{n\rightarrow\infty}\|f_{n}^{\prime}-f_{0}^{\prime}\|_{L^{p}({\mathcal{I}})}\neq 0$ for each $1\leq p\leq\infty$. Consequently, the convergence of the plug-in Doob’s guidance estimator $\nabla\log\widehat{h}_{t}$ used by Uehara et al. (2025b); Ouyang et al. (2024) cannot be guaranteed in general.
@@
To address this issue, Tang and Xu (2024) estimates the Doob’s $h$-function and its gradient separately through a martingale approach. In contrast, this work proposes an approach that simultaneously estimate the Doob’s $h$-function and its gradient.
Doob’s matching with gradient penalty
@@
To simultaneously estimate the Doob’s $h$-function and its gradient, we adopt gradient-penalized regression (Drucker and Le Cun, 1991, 1992; Ding et al., 2025). The population risk is defined by incorporating an additional Sobolev regularization to (3.16):
@@
(3.19) ${\mathcal{J}}_{t}^{\lambda}(h_{t})$ $\coloneq{\mathcal{J}}_{t}(h_{t})+\overbrace{\lambda{\mathbb{E}}^{{\mathbb{P}}}\big[\|\nabla h_{t}(\mathbf{X}_{t}^{\leftarrow})\|_{2}^{2}\big]}^{\text{gradient penalty}}$
@@
where $\lambda>0$ is a regularization parameter.
By a similar argument as (3.17), we can also introduce an empirical counterpart of (3.19):
@@
where $\mathbf{X}_{0}^{1},\ldots,\mathbf{X}_{0}^{n}$ are independent and identically distributed random variables random variables drawn from the base distribution $p_{0}$, and ${\bm{\varepsilon}}_{1},\ldots,{\bm{\varepsilon}}_{n}$ are independent standard Gaussian random variables. Then we have a gradient penalized empirical risk minimizer:
@@
Remark 3.5.
In Section 4.3, we will demonstrate that not only does the gradient-penalized estimator $\widehat{h}_{t}^{\lambda}$ in (3.21) converges to the Doob’s $h$-function $h_{t}^{*}$, but its gradient $\nabla\widehat{h}_{t}^{\lambda}$ also approaches $h_{t}^{*}$, provided that the regularization parameter $\lambda$ and the hypothesis class ${\mathscr{H}}_{t}$ are chosen appropriately.
Doob’s guidance estimation
The Doob’s matching with gradient penalty (3.21) yields a valid estimator of the Doob’s guidance:
@@
This plug-in Doob’s guidance estimator $\widehat{\mathbf{g}}_{t}^{\lambda}$ is characterized by the gradient-penalized regressor $\widehat{h}_{t}^{\lambda}$ in (3.21), thus termed the regressor-guidance.
3.4  A summary of procedure

### v2

@@ 3.2  Controllable diffusion models with Doob’s $h$-transform
In this subsection, we achieve inference-time alignment by incorporating guidance into the reference diffusion models via Doob’s $h$-transform (Rogers and Williams, 2000; Särkkä and Solin, 2019; Heng et al., 2024; Chewi, 2025). We begin by constructing a target path measure $\mathbb{Q}$} on the filtered probability space $(\Omega,{\mathcal{F}},{\mathbb{F}})$. We assume that $\mathbb{Q}$} is absolutely continuous with respect to the reference path measure $\mathbb{P}$} defined in Section (2.2). By the Radon-Nikodym theorem, this relationship is characterized by the existence of an $\mathbb{F}$}-adapted process $(L_{t})_{0\leq t\leq T}$ with $L_{t}\geq 0$, such that for every $t\in[0,T]$:
@@
The process $L_{t}$ is known as the Radon-Nikodym derivative process, representing the likelihood ratio between the target and reference measures conditioned on the filtration ${\mathcal{F}}_{t}$.
@@
The rest of our derivation proceeds in two steps. First, we characterize the stochastic dynamics of $\mathbf{X}_{t}^{\leftarrow}$ under the target measure $\mathbb{Q}$}. Second, by invoking the weak uniqueness of solutions to stochastic differential equations (Øksendal, 2003, Lemma 5.3.1), we construct a controllable diffusion process under the reference path measure $\mathbb{P}$} whose terminal distribution coincides with the target tilted distribution $q_{0}$.
@@
where $h^{*}:[0,T]\times{\mathbb{R}}^{d}\to{\mathbb{R}}$, referred to as Doob’s $h$-function, is defined as the conditional expectation of the terminal weight:
@@
The proof of Proposition 3.1 is provided in Appendix A. With the dynamics of $\log L_{t}$ established, we derive the stochastic dynamics of $\mathbf{X}_{t}^{\leftarrow}$ under the target measure $\mathbb{Q}$} via Girsanov’s theorem (Øksendal, 2003, Theorem 8.6.6).
@@
Let the reference process $\mathbf{X}_{t}^{\leftarrow}$ satisfy the SDE (2.4) under the path measure $\mathbb{P}$}, and let $\mathbb{Q}$} be the target path measure defined by (3.9). Assume that the Novikov condition holds:
@@
where $\mathbf{B}_{t}$ is a Brownian motion under $\mathbb{P}$}. Then $\widetilde{\mathbf{B}}_{t}$ is a standard Brownian motion under $\mathbb{Q}$}. Further, under the path measure $\mathbb{Q}$}, the reference process $\mathbf{X}_{t}^{\leftarrow}$ in (2.4) evolves according to:
@@
The proof is deferred to Appendix A. By the construction in (3.9), the law of $\mathbf{X}_{T}^{\leftarrow}$ under $\mathbb{Q}$} coincides with the target tilted distribution $q_{0}$.
@@
Controllable diffusion process under the reference measure
@@
To address this, we construct a surrogate process $\mathbf{Z}_{t}^{\leftarrow}$ on the reference probability space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$ that adopts the drift derived for $\mathbb{Q}$}:
@@
4  Variationally Stable Doob’s Matching
We have thus far established a controllable diffusion process (3.14) capable of generating samples from the target tilted distribution $q_{0}$. However, the Doob’s $h$-guidance, required by (3.14), remains intractable. This subsection proposes a variationally stable Doob’s matching method to address the estimation of the Doob’s guidance.
4.1  Vanilla least-squares regression for Doob’s matching
For any $t\in(0,T)$, the Doob’s $h$-function $h_{t}^{*}\coloneq h^{*}(t,\cdot)$ defined as (3.10) is the unique minimizer of the following implicit Doob’s matching objective:
@@ 4.1  Vanilla least-squares regression for Doob’s matching
$=\mathbb{E}_{\mathbf{X}_{0}\sim p_{0}}\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})}\big[\|h_{t}(\mu_{T-t}\mathbf{X}_{0}+\sigma_{T-t}\bm{\epsilon})-w(\mathbf{X}_{0})\|_{2}^{2}\big],$
where $w:\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ is the known weight function in (3.1). The following proposition justifies the use of $\mathcal{J}_{t}$ as a surrogate for the explicit $L^{2}$-distance.
@@
For every $t\in(0,T)$, the Doob’s $h$-function $h_{t}^{*}$ in (3.10) minimizes the implicit Doob’s matching objective (4.1). Further,
@@
The proof of Proposition 4.1 is provided in Appendix B.
4.2  Limitations of vanilla regression
@@
Crucially, computing the Doob’s guidance $\nabla\log h_{t}^{*}$ in (3.14) requires estimating not only the function $h_{t}^{*}$ itself but also its gradient $\nabla h_{t}^{*}$, as the guidance is given by:
$\nabla\log h_{t}^{*}(\mathbf{x})=\frac{\nabla h_{t}^{*}(\mathbf{x})}{h_{t}^{*}(\mathbf{x})},\quad\mathbf{x}\in{\mathbb{R}}^{d}.$
According to Proposition 4.1, the objective $\mathcal{J}_{t}$ is only coercive with respect to the $L^{2}(p_{T-t})$-norm. Specifically, for any $h_{t}\in L^{2}(p_{T-t})$,
(4.2) ${\mathcal{J}}_{t}(h_{t})-{\mathcal{J}}_{t}(h_{t}^{*})=\|h_{t}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}.$
However, even if $h_{t}$ is close to $h_{t}^{*}$ in the $L^{2}$ sense, the gradient $\nabla h_{t}$ may remain highly oscillatory. This leads to an unstable plug-in estimator for the Doob’s guidance, a difficulty noted in prior works such as Tang and Xu (2024, Section 3.2.1) and Mou (2025, Section 3.2.2). Similar issues arise in related contexts, including classifier guidance (Dhariwal and Nichol, 2021), Monte Carlo regression (Uehara et al., 2025b, Section 2.2), and Ouyang et al. (2024).
To illustrate this fundamental limitation, consider the sequence of functions $f_{n}:{\mathcal{I}}\to{\mathbb{R}}$ defined by $x\mapsto n^{-1}\sin(nx)$, alongside the zero function $f_{0}\equiv 0$, where ${\mathcal{I}}:=[0,2\pi]$. While $\|f_{n}-f_{0}\|_{L^{2}({\mathcal{I}})}\to 0$ as $n\to\infty$, the distance between their derivatives does not vanish, i.e., $\lim_{n\to\infty}\|f_{n}^{\prime}-f_{0}^{\prime}\|_{L^{2}({\mathcal{I}})}\neq 0$. As a result, the convergence of the plug-in Doob’s guidance estimator derived from vanilla regression (4.1), as utilized in Uehara et al. (2025b); Ouyang et al. (2024), cannot be guaranteed in general.
@@
To mitigate this, Tang and Xu (2024) estimate Doob’s $h$-function and its gradient separately via a martingale approach. In contrast, in the remainder of this work, we propose an approach to simultaneously estimate both the function and its gradient.
4.3  Variationally stable Doob’s matching
@@
To simultaneously estimate the Doob’s $h$-function and its gradient, we adopt a gradient-regularized regression (Drucker and Le Cun, 1991, 1992; Ding et al., 2025b). The population risk is defined by incorporating an additional Sobolev regularization to (4.1):
@@
(4.3) $h_{t}^{\lambda}=\operatorname*{arg\,min}_{h_{t}:{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}}{\mathcal{J}}_{t}^{\lambda}(h_{t})$ $\coloneq{\mathcal{J}}_{t}(h_{t})+\overbrace{\lambda{\mathbb{E}}^{{\mathbb{P}}}\big[\|\nabla h_{t}(\mathbf{X}_{t}^{\leftarrow})\|_{2}^{2}\big]}^{\text{gradient regularization}}$
@@
where $\lambda>0$ is a regularization parameter. The following results characterize the regularization gap and the variational stability of this formulation.
Proposition 4.2 (Regularization gap).
Let $\lambda>0$, $h_{t}^{*}$ be the Doob’s $h$-function defined as (3.10), and $h_{t}^{\lambda}$ be the minimizer of ${\mathcal{J}}_{t}^{\lambda}$ defined as (4.3). Then $h_{t}^{*}\in H^{2}(p_{T-t})$, and
$\|h_{t}^{\lambda}-h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$ $\leq\lambda^{2}\|\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}^{2},$
$\|\nabla h_{t}^{\lambda}-\nabla h_{t}^{*}\|_{L^{2}(p_{T-t})}^{2}$ $\leq\lambda\|\Delta h_{t}^{*}+\nabla h_{t}^{*}\cdot\nabla\log p_{T-t}\|_{L^{2}(p_{T-t})}^{2}.$
The proof of Proposition 4.2 is provided in Appendix B. This proposition demonstrates that as $\lambda\to 0$, the minimizer $h_{t}^{\lambda}$ of the regularized objective (4.3) converges to $h_{t}^{*}$ in $H^{1}$-norm. More importantly, the objective is variationally stable in the $H^{1}$ sense:
Proposition 4.3 (Variational stability).
Let $\lambda>0$, and $h_{t}^{\lambda}$ be the minimizer of ${\mathcal{J}}_{t}^{\lambda}$ defined as (4.3). Then for any $h_{t}\in H^{1}(p_{T-t})$, we have
$\frac{1}{\max\{\lambda,1\}}\big\{{\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})\big\}\leq\|h_{t}-h_{t}^{\lambda}\|_{H^{1}(p_{T-t})}^{2}\leq\frac{1}{\min\{\lambda,1\}}\big\{{\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})\big\}.$
The proof of Proposition 4.3 is provided in Appendix B. We refer to the regularized objective ${\mathcal{J}}_{t}^{\lambda}$ as variationally stable because the convergence in this objective functional necessitates simultaneous convergence in both the function values and their gradients, i.e., stability in the $H^{1}$ sense. Such variational stability ensures that any candidate function $h_{t}$ achieving a low objective value $\mathcal{J}_{t}^{\lambda}(h_{t})$ is guaranteed to be an approximation of the ground-truth Doob’s $h$-function in both value and gradient.
Comparison between vanilla and gradient-regularized Doob’s matching
The distinction between vanilla Doob’s matching (4.1) and the proposed gradient-regularized Doob’s matching (4.3) is fundamental for stable diffusion guidance estimation. While vanilla regression guarantees convergence in the $L^{2}$-norm; see (4.2), it can be unstable in the $H^{1}$ sense. In contrast, the proposed objective ${\mathcal{J}}_{t}^{\lambda}$ is coercive with respect to the $H^{1}$ norm; that is, for a fixed $\lambda>0$,
${\mathcal{J}}_{t}^{\lambda}(h_{t})-{\mathcal{J}}_{t}^{\lambda}(h_{t}^{\lambda})\simeq\|h_{t}-h_{t}^{\lambda}\|_{H^{1}(p_{T-t})}^{2}.$
This property ensures the simultaneous estimation of Doob’s $h$-function and its gradient, thereby yielding a plug-in estimator for Doob’s guidance with mathematical guarantees.
4.4  Doob’s guidance estimation
Since the expectation in the population risk (4.3) is computationally intractable, we approximate it by empirical risk using independent and identically distributed samples:
@@ 4.4  Doob’s guidance estimation
where the empirical least-squares risk is defined as
(4.5) $\widehat{{\mathcal{J}}}_{t}(h_{t})\coloneq\frac{1}{n}\sum_{i=1}^{n}\|h_{t}(\mu_{T-t}\mathbf{X}_{0}^{i}+\sigma_{T-t}{\bm{\varepsilon}}^{i})-w(\mathbf{X}_{0}^{i})\|_{2}^{2}.$
Here $\mathbf{X}_{0}^{1},\ldots,\mathbf{X}_{0}^{n}$ are independent and identically distributed random variables drawn from the reference distribution $p_{0}$, and ${\bm{\varepsilon}}_{1},\ldots,{\bm{\varepsilon}}_{n}$ are independent standard Gaussian random variables. Then one has a gradient-regularized empirical risk minimizer:
@@
The Doob’s matching with gradient regularization (4.6) yields a valid plug-in estimator of the Doob’s guidance:
@@
4.5  A summary of computing procedure